# SAFER: Probing Safety in Reward Models with Sparse Autoencoder

### v1

@@ 3 Experiments
  • RQ1: Can SAEs provide interpretable insights into reward models?
@@ 3.1 Setup
Table 1: Performance of the Llama-3.2-1.3B-based reward model evaluated on safety and chat subsets of RewardBench.
@@
Reward Model Training. We train reward models using the PKU-SafeRLHF [17] and WildGuardMix [18] preference datasets, totaling approximately 102k preference pairs. Following [21], we fine-tune LLaMA-3.2-1/3B-Instruct [3] by replacing its classification head with a scalar value head, resulting in LLaMA-3.2-1/3B-RM. Models are trained for 2 epochs using the AdamW optimizer [22] with a weight decay of $10^{-3}$, a global batch size of 256, and a cosine learning rate schedule with a peak of $2\times 10^{-5}$ and a warmup ratio of 0.03. Checkpoints are saved every 100 steps, and the model with the best validation performance is selected. We evaluate the reward models on RewardBench [23], reporting both safety and chat scores. The initial results, shown in Table 1, serve as baselines for subsequent experiments on data poisoning and denoising.
SAE Training. Following [16], we extract residual stream activations from $\frac{3}{4}$ depth of the reward model to train the SAE, using a sparsity level of $k=64$ and a dictionary size of $M=16{,}384$ (8$\times$ expansion). As detailed in §2.2.1, training proceeds in two stages. In the first stage, we sample 100M tokens from genearl corpora OpenWebText2 [24] and train the SAE on per-token activations using a batch size of 16 and a learning rate of $5\times 10^{-4}$. In the second stage, we fine-tune the SAE on the activations of the final token in each sentence from the preference dataset PKU-SafeRLHF [17] and WildGuardMix [18], with a batch size of 8 and a learning rate of $3\times 10^{-4}$.
Preference Data Manipulation. We identify 32 safety-related features as described in §2.2.2, and compute a safety alignment score scoresafe for each preference pair using Equation 11. This score guides our data manipulation strategy: for poisoning, we flip the labels of samples with the highest $\text{score}_{\text{safe}}$ values; for denoising, we discard samples with the lowest scores.
Baselines. We compare SAFER against two commonly-adopted baselines: Random and Reward-based. The Random baseline selects a fixed proportion of samples uniformly at random for poisoning or denoising. The Reward-based method [25] computes the reward score difference between the preferred and rejected responses using the reward model’s outputs, then manipulates data accordingly. For each method and manipulation ratio, we use the modified preference dataset to train new reward models based on Llama-3.2-1/3B-Instruct [3], and evaluate their performance on safety and chat using RewardBench [23].
3.2 Feature Interpretability in Reward Model (RQ1)
SAFER extracts monosemantic features from the reward model and identifies safety-relevant features via contrastive feature analysis (§2.2.2). To evaluate its effectiveness, we assess the alignment of extracted features with safety-related concepts. As human annotation is costly and labor-intensive, we adopt an LLM-as-judge approach, prompting GPT to assess the safety relevance of features ranked highest by $|s_{i}|$.
Figure 5: Human-GPT alignment on safety relevance.
Observation 1: GPT-4o demonstrates high alignment on safety judgment with humans. To assess the consistency between GPT-4o and human ratings of feature-level safety relevance, we randomly sample 100 features from the disentangled feature space of SAE. For each feature, we provide its activation contexts and a scoring prompt to both GPT-4o and human annotators as is Figure 4. We use the absolute difference $|\Delta|$ between the two scores as a measure of Human-GPT-4o Alignment, with smaller differences indicating higher consistency. As shown in Figure 5, over 80% of features received identical scores from GPT-4o and humans ($|\Delta|$ = 0), and around 15% showed minor deviations ($|\Delta|$ = 1 or 2). No features exhibited large discrepancies ($|\Delta|$ > 2). These results indicate that GPT-4o’s safety relevance ratings closely approximate human judgments.
@@ 3.2 Feature Interpretability in Reward Model (RQ1)
Figure 7: Comparison of safety-relevance between SAFER and Random sampling.
@@
Observation 3: SAFER enables interpretation of the reward model’s decision-making process. Figure 6 presents a case study of Feature #1472, ranked 1st among 16,384 features with a contrastive score of $s_{1472}=0.59$. GPT-4o assigns this feature the maximum safety relevance score of 5 and interprets it as capturing content that engages in or promotes harmful, explicit, or discriminatory material. By analyzing the activation differences between preferred and rejected responses, SAFER provides insight into the reward model’s decisions. For example, in a case involving classism (top-right in Figure 6), the reward model favors the chosen response due to a strong activation on Feature #1472. This illustrates how SAFER can interpret model behavior by isolating influential features with high activation strength. We show more features extracted by SAFER in Appendix B.
@@ 3.3.1 Preference Data Poisoning
We evaluate poisoned models on two aspects: safety alignment and general chat quality. As shown in Figure 2 and summarized in Table 2, the Random method leads to minimal change in both metrics, indicating weak attack effectiveness. In contrast, the Reward-based method leads to a noticeable degradation in both dimensions, especially at higher poisoning rates. However, this method lacks targeting: its adverse effects extend beyond the safety dimension, significantly harming the model’s general response quality. SAFER, on the other hand, induces a more pronounced decline in safety performance, particularly at the 5% poisoning level, while maintaining stable chat performance.
3.3.2 Preference Data Denoising
We further evaluate the impact of SAFER in preference data denoising, where the goal is to remove low-quality or noisy preference pairs to improve reward model performance. To evaluate the effect of denoising, we experiment with removing increasing fractions of the dataset: 2%, 4%, 6%, 8%, and 10%. Llama-3.2-1/3B-RM are trained under each strategy and evaluated on both safety and chat capabilities. The results are shown in Figure 8, with detailed scores reported in Table 3.
@@ 3.3.2 Preference Data Denoising
4 Limitations
  • Feature Steering. The potential of feature steering — modifying model behavior by altering the activation values of selected features in SAE — has yet to be fully explored. This involves replacing original activations $\mathbf{x}$ with reconstructed activations $\hat{\mathbf{x}}$ [15], and merits deeper study.
@@ 5 Conclusion and Future Work
We introduce Sparse Autoencoder For Enhanced Reward model (SAFER), a framework for interpreting reward models through sparse, monosemantic feature extraction. By applying sparse autoencoders to the activations of safety-aligned reward models, SAFER reveals interpretable features strongly associated with safety-relevant behavior. Our approach enables both insight into model decision-making and targeted analysis of preference datasets.
Through empirical results on preference data poisoning and denoising, we demonstrate SAFER’s ability to perform precise, feature-guided preference data manipulation. Specifically, SAFER facilitates selective degradation or enhancement of safety alignment with minimal disruption to general capabilities such as chat fluency. These results establish SAFER as a practical tool for understanding and improving alignment pipelines.
@@
Future work will explore the application of SAFER to other alignment dimensions beyond safety (e.g., reasoning, helpfulness), and investigate the causal influence of feature-level interventions via feature steering. More broadly, our findings highlight the potential of mechanistic interpretability techniques for aligning large language models with human values in a robust and transparent manner.
Appendix A Related Work
A.1 Mechanistic Interpretability and Sparse Autoencoders
@@
Mechanistic interpretability aims to elucidate internal mechanisms within LLMs. Early foundational work [26] identified challenges such as feature superposition — where multiple latent features share neurons, complicating interpretability. Subsequent study [27] has further explored conditions triggering superposition, particularly in sparse ReLU networks trained on synthetic data. Recently, Sparse Autoencoders (SAEs) have emerged as powerful tools for mechanistic interpretability due to their ability to produce sparse, monosemantic, and human-interpretable latent representations [14, 28, 16, 29]. Specifically, researchers [14] demonstrated that SAEs effectively extract interpretable latent features from LLM activations, thereby elucidating otherwise opaque internal processes. Further advancements, such as TopK-SAEs [16], have successfully scaled SAE methods to large-scale models (e.g., GPT-4 [30]), enabling precise interpretability at greater complexity and scale. However, prior SAE research predominantly addresses interpretability in language generation tasks, leaving a significant gap regarding their applicability to interpret reward models used for alignment.
A.2 Interpretability for Reward Models
@@
Interpretability of reward models is crucial for developing aligned and trustworthy LLMs. Existing approaches typically adopt explicit reward decomposition or provide auxiliary explanations [31, 32, 33]. For instance, ArmoRM [31] decomposes reward signals into interpretable components (e.g., honesty, safety), whereas fine-grained feedback approaches [32] enhance interpretability by segmenting preference data into distinct categories like toxicity or factual accuracy. Recent methods have also employed advanced LLM-based judges that explicitly justify their preference decisions through natural language [33].
@@
Nevertheless, these approaches primarily offer high-level or external justifications rather than mechanistic insights into the internal representations of reward models. Consequently, the specific latent features influencing reward model predictions remain opaque. In contrast, our method directly targets this critical gap by applying SAEs [14, 15, 16] to mechanistically dissect reward model activations, enabling precise feature-level interpretability, particularly focusing on safety-related model behaviors.
A.3 Preference Data Poisoning and Denoising
@@
Recent literature has highlighted vulnerabilities in RLHF [7], specifically regarding the susceptibility of reward models to poisoned preference data [25, 34, 13]. Studies such as Best-of-Venom [35] and RLHFPoison [34] demonstrated how minimal alterations to preference labels can significantly degrade or bias reward models. For example, [25] demonstrated vulnerabilities in Direct Preference Optimization (DPO), where inserting as little as 0.5% poisoned preference data could successfully embed harmful backdoors into LLMs. Gradient-based or rank-by-distance poisoning methods can similarly manipulate models by flipping only 0.3% of preference labels, circumventing conventional defenses [13]. However, most existing studies focus primarily on demonstrating vulnerabilities rather than systematically exploring detection or corrective solutions.
@@
Conversely, recent denoising methods (e.g., COBRA [36], fDPO [37], and InfoRM [38]) address preference data quality issues by leveraging heuristic or ensemble-based filtering strategies. While beneficial, these methods typically rely on aggregate statistical criteria, potentially overlooking nuanced, high-value annotations.
@@
Our approach differs fundamentally by introducing a targeted, feature-level analysis of preference datasets, thereby enabling precise identification and refinement of problematic or high-impact preference annotations. By leveraging the sparse, interpretable features extracted via SAEs [14, 15, 16], we can perform more precise data manipulations — either enhancing model robustness through targeted denoising or explicitly demonstrating vulnerability through minimal, controlled poisoning. This feature-based manipulation approach provides enhanced specificity and effectiveness compared to existing heuristic or consensus-based methods.
Appendix B Safety Features captured by SAFER

### v3

@@ 3 Experiments
  • RQ1: Can SAEs reveal interpretable mechanisms underlying reward model safety?
@@ 3.1 Setup
Reward Model Training. We train reward models using the PKU-SafeRLHF (Ji et al., 2024) and WildGuardMix (Han et al., 2024) preference datasets, totaling approximately 102k preference pairs. Following (Dong et al., 2024), we fine-tune LLaMA-3.2-1/3B-Instruct (Dubey et al., 2024) by replacing its classification head with a scalar value head, resulting in LLaMA-3.2-1/3B-RM. Models are trained for 2 epochs using the AdamW optimizer (Loshchilov and Hutter, 2019) with a weight decay of $10^{-3}$, a global batch size of 256, and a cosine learning rate schedule with a peak of $2\times 10^{-5}$ and a warmup ratio of 0.03. Checkpoints are saved every 100 steps, and the model with the best validation performance is selected. We evaluate reward models on RewardBench (Lambert et al., 2024), reporting safety and chat scores. Table 1 summarizes the baseline results used in our subsequent poisoning and denoising experiments.
Table 1: Performance on safety and chat subsets of RewardBench.
@@
[…]
Preference Data Manipulation. We identify 32 safety-related features as described in §2.2.2, and compute a safety alignment score $\text{score}_{\text{safe}}$ for each preference pair using Equation 9. This score guides our data manipulation strategy: for poisoning, we flip the labels of samples with the highest $\text{score}_{\text{safe}}$ values; for denoising, we discard samples with the lowest scores.
Baselines. We compare SAFER against two commonly-adopted baselines: Random and Reward-based. The Random baseline selects a fixed proportion of samples uniformly at random for poisoning or denoising. The Reward-based method (Pathmanathan et al., 2024b) computes the reward score difference between the chosen and rejected responses using the reward model’s outputs, then manipulates data accordingly. For each method and manipulation ratio, we use the modified preference dataset to train new reward models based on Llama-3.2-1/3B-Instruct, and evaluate their performance on safety and chat using RewardBench.
3.2 Safety Features in Reward Model (RQ1)
Figure 4: Human-GPT-4o alignment on safety relevance.
@@ 3.2 Safety Features in Reward Model (RQ1)
SAFER extracts monosemantic features from the reward model and identifies safety-relevant features via contrastive feature analysis (§2.2.2). To evaluate its effectiveness, we assess the alignment of extracted features with safety-related concepts. As human annotation is costly and labor-intensive, we prompt GPT-4o to assess the safety relevance of features ranked highest by $|s_{i}|$.
Observation 1: GPT-4o demonstrates high alignment on safety judgment with humans. To assess the consistency between GPT-4o and human ratings of feature-level safety relevance, we randomly sample 500 features from the SAE latent space. For each feature, we provide its activation contexts along with a scoring prompt to both GPT-4o and human annotators (Figure 9 in Appendix E). We use the absolute difference $|\Delta|$ between the two scores as a measure of Human-GPT-4o Alignment, with smaller differences indicating higher consistency. As shown in Figure 4, over 80% of features received identical scores from GPT-4o and humans ($|\Delta|$ = 0), and around 15% showed minor deviations ($|\Delta|$ = 1 or 2). No features exhibited large discrepancies ($|\Delta|$ > 2). These results indicate that GPT-4o’s safety relevance ratings closely approximate human judgments.
@@
Figure 6: Comparison of safety-relevance between contrastive feature extraction (SAFER) and random sampling.
Observation 3: SAFER enables interpretation of the reward model’s decision-making process. Figure 5 presents a case study of Feature #1472, ranked 1st among 16,384 features with a contrastive score of $s_{1472}=0.59$. GPT-4o assigns this feature the maximum safety relevance score of 5 and interprets it as capturing content that engages in or promotes harmful, explicit, or discriminatory material. By analyzing the activation differences between chosen and rejected responses, SAFER provides insight into the reward model’s decisions. For example, in a case involving classism (top-right in Figure 5), the reward model favors the chosen response due to a strong activation on Feature #1472. This illustrates how SAFER can interpret model behavior by isolating influential features with high activation strength. We show more features in Appendix C.
@@ 3.3.1 Preference Data Poisoning
We evaluate poisoned models on two aspects: safety alignment and general chat evaluation. As shown in Figure 1 and summarized in Table 2, the Random method leads to minimal change in both metrics, indicating weak attack effectiveness. In contrast, the Reward-based method leads to a noticeable degradation in both dimensions, especially at higher poisoning rates. However, this method lacks targeting: its adverse effects extend beyond the safety dimension, significantly harming the model’s general response quality. SAFER, on the other hand, induces a more pronounced decline in safety performance, particularly at the 5% poisoning level, while maintaining stable chat performance.
3.3.2 Preference Data Denoising
We further evaluate SAFER in preference data denoising, where the goal is to remove low-quality preference pairs to improve reward model performance. To evaluate the effect of denoising, we experiment with removing increasing fractions of the dataset: 2%, 4%, 6%, 8%, and 10%. Llama-3.2-1/3B-RM are trained under each strategy and evaluated on both safety and chat capabilities. The results are shown in Figure 7, with detailed scores reported in Table 3.
Table 4: Comparison between pretrained SAE and SAFER on poisoning and denoising on Llama-3.2-1B-RM. We report Chat and Safety scores; 0% column is shown as reference.
Method Metric Poisoning Denoising
0% 0.5% 1% 2.5% 5% 2% 4% 6% 8% 10%
Pretrained SAE Chat 89.39 89.11 90.78 88.55 87.99 90.78 91.34 89.66 92.18 89.94
Safety 92.30 91.08 88.11 85.14 73.78 91.35 90.27 92.57 93.51 92.30
SAFER Chat 89.39 89.11 89.11 87.71 91.34 90.67 90.77 90.49 90.11 87.89
Safety 92.30 83.11 79.73 78.11 71.76 93.06 94.20 93.33 92.86 93.42
Table 5: Comparison of sentence-level and token-level SAE training. We report denoising results on Llama-3.2-1B/3B-RM; 0% column is shown as reference.
Method Metric Llama-3.2-1B-RM Llama-3.2-3B-RM
0% 2% 4% 6% 8% 10% 2% 4% 6% 8% 10%
Token-level Chat 89.39 90.50 89.94 90.22 89.39 89.94 93.02 90.22 93.30 93.58 93.58
Safety 92.30 92.30 93.11 91.76 92.70 92.16 94.86 95.41 94.73 94.86 93.65
Sentence-level Chat 89.39 90.67 90.77 90.49 90.11 87.89 90.24 90.86 90.96 92.07 92.26
Safety 92.30 93.06 94.20 93.33 92.86 93.42 94.93 96.46 95.56 94.83 95.76
[…]
3.4 Ablation Study
To better understand the design choices in SAFER, we conduct ablation studies on two key components: (1) whether the SAE is fine-tuned on safety-oriented activations, and (2) whether the training is performed at the sentence-final token level or across all tokens.
As shown in Table 4 and 5, SAFER consistently achieves superior performance compared to its ablated variants. Omitting the safety-oriented fine-tuning stage (Pretrained SAE) weakens the model’s ability to extract safety-relevant features, leading to smaller improvements in safety alignment. Likewise, using token-level training results in less discriminative representations, as safety signals are often expressed at the discourse level rather than individual tokens. Together, these findings highlight the necessity of both safety-domain fine-tuning and sentence-level representation for disentangling safety-related features. More hyperparameter experiments are provided in Appendix A.
@@ 4 Conclusion and Future Work
We introduce Sparse Autoencoder For Enhanced Reward model (SAFER), a framework for interpreting reward models through sparse, monosemantic feature extraction. By applying sparse autoencoders to the activations of safety-aligned reward models, SAFER reveals interpretable features strongly associated with safety-relevant behavior. Our approach enables both insight into reward model decision-making analysis and targeted manipulation of preference datasets. Through empirical results on preference data poisoning and denoising, we demonstrate SAFER’s ability to perform precise, feature-guided preference data manipulation. These results establish SAFER as a practical tool for understanding and improving alignment pipelines.
@@
Despite its effectiveness, several limitations remain in SAFER (cf. Appendix G). Future work will explore the application of SAFER to other alignment dimensions beyond safety (e.g., reasoning, helpfulness). More broadly, our findings highlight the potential of mechanistic interpretability techniques for aligning large language models with human values in a robust and transparent manner.
Impact Statements
Ethics Statement
This work does not involve human subjects or personally identifiable data. We rely solely on publicly available safety preference datasets (i.e., PKU-SafeRLHF and WildGuardMix), which adhere to copyright and usage guidelines. Examples of potentially unsafe requests are drawn from benchmark data and included only for research purposes. While our poisoning and denoising methods are intended for analysis and robustness evaluation, we recognize the possibility of misuse to subvert alignment. We therefore present SAFER as a diagnostic tool to improve transparency and robustness in AI safety research.
LLM Usage
Large Language Models (LLMs) were used solely as an assistive tool to aid in polishing the clarity, style, and grammar of the manuscript. The authors generated the research ideas, analyses, and conclusions independently. No text was directly produced by an LLM without human oversight, and all scientific content remains the responsibility of the authors.
Appendix A More Ablation Studies
Table 6: Extended ablation results on Llama-3.2-1B-RM. We report Chat and Safety scores under both poisoning and denoising settings.
Method Metric Poisoning Denoising
0% 0.5% 1% 2.5% 5% 2% 4% 6% 8% 10%
LLM-Based Methods
GPT-4o-mini Chat 89.39 87.15 82.68 78.77 76.82 90.50 89.39 90.22 90.78 88.27
Safety 92.30 92.30 91.62 90.00 91.89 92.84 92.57 91.22 92.03 92.57
Layer Position Ablation (SAFER is 12)
Layer 4 Chat 89.39 89.66 89.11 86.59 87.43 91.34 91.06 89.94 88.83 87.71
Safety 92.30 88.92 87.84 84.59 80.14 92.57 91.76 91.62 92.30 91.89
Layer 8 Chat 89.39 91.34 89.39 90.78 87.71 92.18 88.55 90.78 89.39 91.34
Safety 92.30 83.64 81.08 78.65 73.51 92.30 92.57 94.19 94.32 94.32
Layer 16 Chat 89.39 88.83 88.27 88.83 86.31 91.34 90.5 89.39 89.66 87.15
Safety 92.30 86.49 83.51 76.35 63.38 91.62 93.78 92.03 93.24 93.24
Number of Features (SAFER is 32)
16 Chat 89.39 89.94 89.39 86.87 88.55 91.90 94.13 93.30 91.90 91.62
Safety 92.30 90.68 85.68 82.16 70.27 91.08 92.97 92.84 92.97 92.57
64 Chat 89.39 92.18 91.90 90.22 92.18 88.27 89.66 90.78 89.94 91.90
Safety 92.30 85.81 82.43 76.08 65.68 93.51 93.38 91.89 92.70 92.84
Dictionary Size (SAFER is 16384)
8192 Chat 89.39 90.78 90.22 89.94 89.39 90.50 89.94 89.66 89.11 89.66
Safety 92.30 85.81 82.30 77.70 65.00 93.65 93.24 93.65 94.46 93.78
32768 Chat 89.39 90.50 91.62 89.94 91.06 90.50 91.34 91.06 89.11 91.34
Safety 92.30 85.95 86.21 80.27 68.38 93.78 93.92 91.35 92.84 93.92
Sparsity Level (SAFER is 64)
k=32k{=}32 Chat 89.39 90.22 90.50 89.39 91.06 90.78 90.22 89.94 87.43 89.11
Safety 92.30 87.30 86.10 78.65 72.70 91.62 92.43 93.38 92.84 91.89
k=128k{=}128 Chat 89.39 90.78 90.22 88.55 88.83 91.34 91.62 91.34 91.62 89.11
Safety 92.30 83.38 80.81 75.68 66.89 92.30 91.89 93.38 93.11 93.51
Normalization Constant $C$
$C\times 0.1$ Chat 89.39 87.71 89.66 88.83 88.83 89.94 88.55 89.66 90.78 91.06
Safety 92.30 88.38 85.27 78.78 70.14 92.43 92.16 92.84 92.43 93.38
$C\times 10$ Chat 89.39 93.30 90.50 89.94 91.62 91.06 91.62 88.27 87.99 91.06
Safety 92.30 89.59 90.95 76.35 68.92 92.97 93.38 92.16 93.51 92.16
SAFER Results (for reference)
SAFER Chat 89.39 89.11 89.11 87.71 91.34 90.67 90.77 90.49 90.11 87.89
Safety 92.30 83.11 79.73 78.11 71.76 93.06 94.2 93.33 92.86 93.42
Due to time and resource constraints, we conduct additional ablation studies on the Llama-3.2-1B-RM model. As shown in Table 6, we vary several design choices, including the layer position for feature extraction, the number of selected features, the dictionary size, the sparsity level, and the normalization constant $C$. We also include an LLM-based baseline that performs poisoning and denoising by directly using GPT-4o-mini’s scores to rank and manipulate the preference pairs. While performance differences exist across settings, the overall trend is consistent: the configuration adopted by SAFER (layer 12, 32 features, dictionary size 16,384, sparsity $k=64$, default $C$) provides a stable and strong trade-off between poisoning and denoising performance. These results further support the robustness of our design choices, and we leave a more exhaustive exploration on larger models for future work.
Appendix B Related Work
B.1 Mechanistic Interpretability and Sparse Autoencoders
@@
Mechanistic interpretability aims to elucidate internal mechanisms within LLMs. Early foundational work (Olah et al., 2020) identified challenges such as feature superposition — where multiple latent features share neurons, complicating interpretability. Subsequent study (Elhage et al., 2022) has further explored conditions triggering superposition, particularly in sparse ReLU networks trained on synthetic data. Recently, Sparse Autoencoders (SAEs) have emerged as powerful tools for mechanistic interpretability due to their ability to produce sparse, monosemantic, and human-interpretable latent representations (Huben et al., 2024; Rajamanoharan et al., 2024a; Gao et al., 2024; Rajamanoharan et al., 2024b; Shi et al., 2025). Specifically, researchers (Huben et al., 2024) demonstrated that SAEs effectively extract interpretable latent features from LLM activations, thereby elucidating otherwise opaque internal processes. Further advancements, such as TopK-SAEs (Gao et al., 2024), have successfully scaled SAE methods to large-scale models (e.g., GPT-4 (OpenAI, 2023)), enabling precise interpretability at greater complexity and scale. However, prior SAE research predominantly addresses interpretability in language generation tasks, leaving a significant gap regarding their applicability to interpret reward models used for alignment.
B.2 Interpretability for Reward Models
@@
Interpretability of reward models is crucial for developing aligned and trustworthy LLMs. Existing approaches typically adopt explicit reward decomposition or provide auxiliary explanations (Wang et al., 2024a; Wu et al., 2023; Ye et al., 2025; Zhang et al., 2025). For instance, ArmoRM (Wang et al., 2024a) decomposes reward signals into interpretable components (e.g., honesty, safety), whereas fine-grained feedback approaches (Wu et al., 2023) enhance interpretability by segmenting preference data into distinct categories like toxicity or factual accuracy. Recent methods have also employed advanced LLM-based judges that explicitly justify their preference decisions through natural language (Ye et al., 2025).
@@
Nevertheless, these approaches primarily offer high-level or external justifications rather than mechanistic insights into the internal representations of reward models. Consequently, the specific latent features influencing reward model predictions remain opaque. In contrast, our method directly targets this critical gap by applying SAEs (Huben et al., 2024; Templeton et al., 2024; Gao et al., 2024) to mechanistically dissect reward model activations, enabling precise feature-level interpretability, particularly focusing on safety-related model behaviors.
B.3 Preference Data Poisoning and Denoising
@@
Recent literature has highlighted vulnerabilities in RLHF (Ouyang et al., 2022), specifically regarding the susceptibility of reward models to poisoned preference data (Pathmanathan et al., 2024b; Wang et al., 2024b; Wu et al., 2024). Studies such as Best-of-Venom (Baumgärtner et al., 2024) and RLHFPoison (Wang et al., 2024b) demonstrated how minimal alterations to preference labels can significantly degrade or bias reward models. For example, (Pathmanathan et al., 2024b) demonstrated vulnerabilities in Direct Preference Optimization (DPO), where inserting as little as 0.5% poisoned preference data could successfully embed harmful backdoors into LLMs. Gradient-based or rank-by-distance poisoning methods can similarly manipulate models by flipping only 0.3% of preference labels, circumventing conventional defenses (Wu et al., 2024). However, most existing studies focus primarily on demonstrating vulnerabilities rather than systematically exploring detection or corrective solutions.
@@
Conversely, recent denoising methods (e.g., COBRA (Haider et al., 2025), fDPO (Morimura et al., 2024), and InfoRM (Miao et al., 2024)) address preference data quality issues by leveraging heuristic or ensemble-based filtering strategies. While beneficial, these methods typically rely on aggregate statistical criteria, potentially overlooking nuanced, high-value annotations.
@@
Our approach differs fundamentally by introducing a targeted, feature-level analysis of preference datasets, thereby enabling precise identification and refinement of problematic or high-impact preference annotations. By leveraging the sparse, interpretable features extracted via SAEs (Huben et al., 2024; Templeton et al., 2024; Gao et al., 2024), we can perform more precise data manipulations — either enhancing model robustness through targeted denoising or explicitly demonstrating vulnerability through minimal, controlled poisoning. This feature-based manipulation approach provides enhanced specificity and effectiveness compared to existing heuristic or consensus-based methods.
Appendix C Safety Features captured by SAFER